OpenStudy (anonymous):

how do you solve x2+8x by completing the square

OpenStudy (jhannybean):

you have the formula \(\large ax^2+bx+c\) and you're given \[\large x^2+8x\]

OpenStudy (zzr0ck3r):

set = to 0 x^2 +8x = 0 take half the second term (x+4)^2 = 4^2 = 16 (x+4)^2-16

OpenStudy (zzr0ck3r):

x^2+8x = (x+4)^2-16

OpenStudy (zzr0ck3r):

I don't know why one would do such a thing...

OpenStudy (zzr0ck3r):

factor out the x and get the same solution...

OpenStudy (jhannybean):